OpenStudy (anonymous):

Solve: x - 5 = -10 Answer -15 15 5 -5

OpenStudy (hihi67):

-5

OpenStudy (hihi67):

sorry, thats wrong, -10

OpenStudy (hihi67):

its definately -10!!!!!!

OpenStudy (anonymous):

its -5

OpenStudy (anonymous):

It's -5 x-5=-10 x=-10+5=-5